# Tight Bounds for Sparsifying Random CSPs

### v1

@@ 4.2 Proof of Upper Bound
Now, let $t\in\{0,1\}^{r}$ be extreme for $R$. We let $\mathcal{S}_{t}$ be the set of all $S\in\binom{[r]}{c}$ such that $t\oplus e_{S}\in R$. For an index $i\in[r]$, we say that $r$ is covered by $t$ if there exists a set $S\in\mathcal{S}_{t}$ such that $i\in S$, and otherwise we say that $i$ is irrelevant for $t$.
@@
Define $\bar{t}:=t\oplus e_{T}$. Note that $\bar{t}$ has distance $c+1$ from $t^{\prime}$. Thus, to avoid a copy of $\mathbf{AND}_{c+1}$ in $R$, there must be some $t^{\prime\prime}\in R$ between $\bar{t}$ and $t^{\prime}$. In other words, there exists $S^{\prime}\subsetneq S\cup\{i_{0}\}$ such that $\bar{t}\oplus e_{S^{\prime}}\in R$. We now have a few cases based on the composition of $S^{\prime}$.
If $S^{\prime}=S$, then $\bar{t}\oplus e_{S^{\prime}}=t\oplus e_{S\cup T}\in R$, as desired. Otherwise, $S^{\prime}\cap S\subsetneq S$. In particular, consider $\bar{t}\oplus e_{S^{\prime}}=t\oplus e_{S^{\prime}\oplus T}\in R$. Note that if $|S^{\prime}\oplus T|<c$, we contradict the fact that $t$ is extreme. If $|S^{\prime}\oplus T|=c$, then $S^{\prime}\oplus T\in\mathcal{S}_{t}$. But, since $S^{\prime}\oplus T\neq S$, there must be some irrelevant coordinate in $S^{\prime}\oplus T$. This contradicts the fact that the coordinate is irrelevant. Finally, assume that $|S^{\prime}\oplus T|>c$. Then, to avoid a copy of $\mathbf{AND}_{c+1}$ in $R$, there must be $S^{\prime\prime}\subseteq S^{\prime}\oplus T$ of size $c$ with $t\oplus e_{S^{\prime\prime}}\in R$, so $S^{\prime\prime}\in\mathcal{S}_{t}$. Again, $S^{\prime\prime}\cap S\subseteq S^{\prime}\cap S\subsetneq S$, so $S^{\prime\prime}$ must contain an irrelevant coordinate, a contradiction.
@@ 5.2 Proof of the Upper Bound
Next, we define $\bar{t}$ as $\bar{t}_{i}=0$ for $i\in\bar{T}$, and otherwise $\bar{t}_{i}=s_{i}$. In this sense, we have that $\mathrm{Ham}(\bar{t},t^{\prime})=c+1$. So, in order to ensure that there is no $\mathbf{AND}_{c+1}$ in $\mathbf{R}$, it must be the case that there is some $t^{\prime\prime}\in R$ which is between $\bar{t}$ and $t^{\prime}$ (here, we use between to mean that $t^{\prime\prime}_{i}\in\{t^{\prime}_{i},\bar{t}_{i}\}$). Now, we let $S^{\prime}\subsetneq\operatorname{supp}(t)\cup\{i_{0}\}$ denote the coordinates where $\bar{t}$ and $t^{\prime\prime}$ disagree. We have a few cases depending on $S$.
@@ 6.1 VCSP Notation
A primary (but not only) factor governing the sparsifiability of $C$ is the presence of $\mathbf{AND}_{k}$, which we define below for $r$-sorted valued relations.
@@
We say an $r$-sorted valued relation $\mathbf{R}:D_{1}\times\cdots\times D_{r}\to\mathbb{Z}^{\geq 0}$ has a restriction to $\mathbf{AND}_{k}$ if there exist sets $E_{1}\subseteq D_{1},\ldots,E_{r}\subseteq D_{r}$ such that $1\leq|E_{i}|\leq 2$, exactly $k$ of the $E_{i}$’s are of size $2$, and all together, $|\operatorname{supp}\mathbf{R}\cap E_{1}\times E_{2}\dots\times E_{r}|=1$.
However, unlike Theorem 5.2, the largest $\mathbf{AND}_{k}$ in $\mathbf{R}$ does not suffice to understand the size of the sparsifier. To see why, consider the following example. Let $D=\{0,1\}$ and $r=2$ and let $\mathbf{R}$ be
@@
One can show that $\mathbf{R}$ has a restriction to $\mathbf{AND}_{1}$ but not $\mathbf{AND}_{2}$. Thus, one might expect that the sparsifiability of the complete $r$-partite instance to be on the order of $\widetilde{\Theta}_{\varepsilon}(n)$. However, for reasons similar to that of Theorem 3.1, Item 4, any sparsifier of $\mathbf{R}$ must have size at least $\Omega_{\varepsilon}(n^{2})$. In short, for every $i\in V_{1}$, we need to preserve at least $\Omega_{\varepsilon}(n)$ edges of the form $(i,j)$ with $j\in V_{2}$ in order to correctly estimate the total contribution of $\mathbf{R}(10)$ and $\mathbf{R}(11)$.
@@
6.2 Uniform Marginals and Non-uniform $\mathbf{AND}$
For succinctness, we define $\vec{D}^{r}:=D_{1}\times\cdots\times D_{r}$. Given a valued relation $\mathbf{R}:\vec{D}^{r}\to\mathbb{Z}^{\geq 0}$ and a relation $R_{0}\subseteq\vec{D}^{r}$, we say that $\mathbf{R}\upharpoonright R_{0}$ is uniform if for all $t\in R_{0}$, $\mathbf{R}(t)$ has the same value. This motivates the definition of a generalized $\mathbf{AND}_{k}$.
@@ 6.2 Uniform Marginals and Non-uniform $\mathbf{AND}$
We say that an $r$-sorted valued relation $\mathbf{R}:\vec{D}^{r}\to\mathbb{Z}^{\geq 0}$ has a restriction to a generalized $\mathbf{AND}_{k}$ if there exists $E_{1}\subseteq D_{1},\ldots,E_{r}\subseteq D_{r}$ as well as a set of $k$ distinguished coordinates $S\subseteq[r]$ with the following properties.
@@
We say that $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a uniform $\mathbf{AND}_{k}$ if every nonzero output has the same value. Otherwise, we say that $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a non-uniform $\mathbf{AND}_{k}$.
Note that $\mathbf{R}$ has a restriction to at least one generalized $\mathbf{AND}_{k}$ if and only if $\mathbf{R}$ has a restriction to $\mathbf{AND}_{k}$ (as in Definition 6.1).
@@
In this case, $\mathbf{R}\upharpoonright D^{2}$ is a non-uniform $\mathbf{AND}_{1}$ because the value of the first coordinate suffices to distinguish between zero and non-zero but the second coordinate still affects the value of the output. We can now state our classification.
@@
Let $C$ be the complete, $r$-partite instance of an $r$-sorted valued relation $\mathbf{R}:\vec{D}^{r}\to\mathbb{Z}^{\geq 0}$ with $n$ variables in each part, and let $c=\max\{k:\mathbf{R}\text{ has a restriction to }\mathbf{AND}_{k}\}$. If every generalized $\mathbf{AND}_{c}$ of $\mathbf{R}$ is uniform, let $\widehat{c}=c$. Otherwise, let $\widehat{c}=c+1$. Then:
@@ 6.3 Proof of the Lower Bound
If $\mathbf{R}$ has a restriction to $\mathbf{AND}_{k}$, then every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega(n^{k})$ constraints.
@@
If $\mathbf{R}$ has a restriction to a non-uniform $\mathbf{AND}_{k}$, then every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega(n^{k+1})$ constraints if $\varepsilon$ is a sufficiently small constant depending on the relation $\mathbf{R}$.
@@
Consider $E_{1}\subseteq D_{1},\ldots,E_{r}\subseteq D_{r}$ as well as a set of $k$ distinguished coordinates $S\subseteq[r]$ such that $\mathbf{R}\upharpoonright\vec{E}^{r}$ which is a non-uniform generalized $\mathbf{AND}_{k}$. Without loss of generality, we may assume that $S=[k]$ and $E_{1}=\cdots=E_{k}=\{0,1\}$ so that if $\mathbf{R}(t)>0$ for some $t\in\vec{E}^{r}$ then $(t_{1},\ldots,t_{k})=(1,\ldots,1)$.
Since $\mathbf{R}\upharpoonright\vec{E}^{r}$ is non-uniform, there exists $s,t\in\vec{E}^{r}$ with $0<\mathbf{R}(s)<\mathbf{R}(t)$. Since $\vec{E}^{r}$ is a product set, we can construct a chain of tuples $s^{0}=s,\ldots,s^{\ell}=t\in\vec{E}^{r}$ such that $\operatorname{Ham}(s^{i},s^{i+1})=1$ throughout the chain and $s^{i}_{j}=1$ whenever $j\in[k]$. Since $\mathbf{R}(s)\neq\mathbf{R}(t)$, there must be two consecutive terms in this sequence which differ in value. Thus, we may assume without loss of generality that $\operatorname{Ham}(s,t)=1$. Furthermore, we may assume that $s_{k+1}\neq t_{k+1}$ and $\mathbf{R}(s)<\mathbf{R}(t)$.
@@ 6.4 Proof of the Upper Bound
Since $\mathrm{sat}_{\mathbf{R},C}(\psi)>0$, there exists $e_{1}\in D_{1},\ldots,e_{r}\in D_{r}$ such that $(e_{1},\ldots,e_{r})\in\operatorname{supp}\mathbf{R}$ and $a_{i,e_{i}}\geq 1$ for all $i\in[r]$. Of all such choices of $(e_{1},\ldots,e_{r})$, pick the one which maximizes the number of $i\in[r]$ for which $e_{i}=d_{i}$. Now for all $i\in[r]$, let $E_{i}:=\{d_{i},e_{i}\}.$ By the maximality condition on the choice of $(e_{1},\ldots,e_{r})$, we have that $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a copy of $\mathbf{AND}_{k}$ for some $k\leq c$. Therefore,
@@
Moving forward, we assume that $\widehat{c}=c$. That is, every generalized $\mathbf{AND}_{c}$ appearing in $\mathbf{R}$ is uniform. With this assumption, we now proceed to prove a stronger upper bound of $\widetilde{O}_{\mathbf{R},\varepsilon}(n^{c})$ on the size of the sparsifier using the methods of Section 5. More precisely, let $R:=\operatorname{supp}\mathbf{R}$ and note that $c=\max\{k:R\text{ has a restriction to }\mathbf{AND}_{k}\}$. Therefore, we can use many results of Section 5 verbatim.
@@
By Proposition 5.6, we have that $t^{\prime}\in E_{1}\times\cdots\times E_{r}$ satisfies $t^{\prime}\in\operatorname{supp}\mathbf{R}$ if and only if $t^{\prime}_{i}=t_{i}$ whenever $t_{i}\neq 0$. Thus, $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a generalized $\mathbf{AND}_{c}$ (Definition 6.3) since $\operatorname{Ham}(t)=c$. Since we assume that every generalized $\mathbf{AND}_{c}$ is uniform, we must have that $\mathbf{R}(s)=\mathbf{R}(t)$, a contradiction.
Second, assume for the sake of contradiction that $\mathbf{R}(t)<\mathbf{R}_{1}(t)$. Thus, there is $s\sim_{I}t$ such that $\mathbf{R}(s)>\mathbf{R}(t)$. In particular, $s\in\operatorname{supp}\mathbf{R}$. If $t\in\operatorname{supp}\mathbf{R}$ as well, then we can get a contradiction similar to the previous argument, so assume that $t\not\in\operatorname{supp}\mathbf{R}$. Note that $\operatorname{Ham}(s)\geq c+1$ as $s$ and $t$ can only differ in coordinates $i\in[r]$ for which $t_{i}=0$. Since $\mathbf{R}$ lacks $\mathbf{AND}_{c+1}$, there must be some $s^{\prime}\in\operatorname{supp}\mathbf{R}$ with $\operatorname{Ham}(s^{\prime})=c$ and for all $i\in[r]$ either $s^{\prime}_{i}=0$ or $s^{\prime}_{i}=s_{i}$. Now $s^{\prime}\neq t$ as $t\notin\operatorname{supp}\mathbf{R}$, and it follows that $\operatorname{supp}(s^{\prime})\neq\operatorname{supp}(t)$ as well. However, $\operatorname{supp}(s^{\prime})\cup\operatorname{supp}(t)\subseteq\operatorname{supp}(s)$. Therefore there exists an $i\in\operatorname{supp}(s)\setminus\operatorname{supp}(t)$ for which $(i,s^{\prime}_{i})=(i,s_{i})\not\in I$. This contradicts the facts that $s\sim_{I}t$. We can thus conclude $\mathbf{R}(t)=\mathbf{R}_{1}(t)$. ∎
@@
Proof of Theorem 6.4, Item 2.

### v2

@@ 4.2 Proof of Upper Bound
Now, let $t\in\{0,1\}^{r}$ be extreme for $R$. We let $\mathcal{S}_{t}$ be the set of all $S\in\binom{[r]}{c}$ such that $t\oplus e_{S}\in R$. For an index $i\in[r]$, we say that $i$ is covered by $t$ if there exists a set $S\in\mathcal{S}_{t}$ such that $i\in S$, and otherwise we say that $i$ is irrelevant for $t$.
@@
Define $\bar{t}:=t\oplus e_{T}$. First, we observe that $\bar{t}\notin R$. Indeed, if $\bar{t}\in R$ and $|T|<c$, then this violates the extremality of $t$. If $|T|=c$, then this violates the irrelevance of the coordinates in $T$. Otherwise, if $|T|>c$, then $t\notin R$, but $\bar{t}:=t\oplus e_{T}\in R$, and so there must be some tuple strictly between $t$ and $\bar{t}$ which is in $R$ to avoid a copy of $\mathbf{AND}_{c+1}$. But this yields some new tuple $\bar{t}^{\prime}=t\oplus e_{T^{\prime\prime}}\in R$, for $T^{\prime\prime}\subsetneq T$, and so we can repeat this argument until we either contradict extremality of $T$ or irrelvance of coordinates in $T$.
Now, we know that $\bar{t}$ has distance $c+1$ from $t^{\prime}$ and that $\bar{t}\notin R$. Thus, to avoid a copy of $\mathbf{AND}_{c+1}$ in $R$, there must be some $t^{\prime\prime}\in R$ between $\bar{t}$ and $t^{\prime}$. In other words, there exists $S^{\prime}\subsetneq S\cup\{i_{0}\}$ such that $\bar{t}\oplus e_{S^{\prime}}\in R$. We now have a few cases based on the composition of $S^{\prime}$.
If $S^{\prime}=S$, then $\bar{t}\oplus e_{S^{\prime}}=t\oplus e_{S\cup T}\in R$, as desired. Otherwise, $S^{\prime}\cap S\subsetneq S$. In particular, consider $\bar{t}\oplus e_{S^{\prime}}=t\oplus e_{S^{\prime}\oplus T}\in R$. Note that if $|S^{\prime}\oplus T|<c$, we contradict the fact that $t$ is extreme. If $|S^{\prime}\oplus T|=c$, then $S^{\prime}\oplus T\in\mathcal{S}_{t}$. But, since $S^{\prime}\oplus T\neq S$, there must be some irrelevant coordinate in $S^{\prime}\oplus T$. This contradicts the fact that the coordinate is irrelevant. Finally, assume that $|S^{\prime}\oplus T|>c$. Then, to avoid a copy of and $\and_{c+1}$ in $R$, there must be $S^{\prime\prime}\subseteq S^{\prime}\oplus T$ of size $c$ with $t\oplus e_{S^{\prime\prime}}\in R$, so $S^{\prime\prime}\in\mathcal{S}_{t}$. Again, $S^{\prime\prime}\cap S\subseteq S^{\prime}\cap S\subsetneq S$, so $S^{\prime\prime}$ must contain an irrelevant coordinate, a contradiction.
@@ 5.2 Proof of the Upper Bound
Next, we define $\bar{t}$ as $\bar{t}_{i}=0$ for $i\in\bar{T}$, and otherwise $\bar{t}_{i}=s_{i}$. As in Proposition 4.6, we first claim that $\bar{t}\notin\mathbf{R}$. Indeed, if $\bar{t}\in\mathbf{R}$, we have a few cases: (1) if $|T|<c$, then this violates the extremality of $0^{r}$, (2) if $|T|=c$, then this violates the irrelevance of the coordinate, value pairs in $(i,\bar{t}_{i}):i\in T$, or (3) if $|T|>c$, because $0^{r}\notin\mathbf{R}$ and $\bar{t}\in\mathbf{R}$, then in order to not violate the maximum arity $\mathbf{AND}$ being $c$, there must be a tuple $\bar{t}^{\prime}$ strictly between $0^{r}$ and $\bar{t}$ such that $\bar{t}^{\prime}\in\mathbf{R}$. Importantly, $\bar{t}^{\prime}$ is only non-zero on irrelevant coordinates, and so we repeat the above argument with cases (1) and (2), eventually reaching a contradiction. Thus, $\bar{t}\notin\mathbf{R}$.
Now, we have that $\mathrm{Ham}(\bar{t},t^{\prime})=c+1$. So, because $\bar{t}\notin\mathbf{R}$, in order to ensure that there is no $\mathbf{AND}_{c+1}$ in $\mathbf{R}$, it must be the case that there is some $t^{\prime\prime}\in R$ which is between $\bar{t}$ and $t^{\prime}$ (here, we use between to mean that $t^{\prime\prime}_{i}\in\{t^{\prime}_{i},\bar{t}_{i}\}$). Now, we let $S^{\prime}\subsetneq\operatorname{supp}(t)\cup\{i_{0}\}$ denote the coordinates where $\bar{t}$ and $t^{\prime\prime}$ disagree. We have a few cases depending on $S$.
@@ 6.1 VCSP Notation
A primary (but not only) factor governing the sparsifiability of $C$ is the presence of and $\and_{k}$, which we define below for $r$-sorted valued relations.
@@
We say an $r$-sorted valued relation $\mathbf{R}:D_{1}\times\cdots\times D_{r}\to\mathbb{Z}^{\geq 0}$ has a restriction to and $\and_{k}$ if there exist sets $E_{1}\subseteq D_{1},\ldots,E_{r}\subseteq D_{r}$ such that $1\leq|E_{i}|\leq 2$, exactly $k$ of the $E_{i}$’s are of size $2$, and all together, $|\operatorname{supp}\mathbf{R}\cap E_{1}\times E_{2}\dots\times E_{r}|=1$.
However, unlike Theorem 5.2, the largest and $\and_{k}$ in $\mathbf{R}$ does not suffice to understand the size of the sparsifier. To see why, consider the following example. Let $D=\{0,1\}$ and $r=2$ and let $\mathbf{R}$ be
@@
One can show that $\mathbf{R}$ has a restriction to and $\and_{1}$ but not and $\and_{2}$. Thus, one might expect that the sparsifiability of the complete $r$-partite instance to be on the order of $\widetilde{\Theta}_{\varepsilon}(n)$. However, for reasons similar to that of Theorem 3.1, Item 4, any sparsifier of $\mathbf{R}$ must have size at least $\Omega_{\varepsilon}(n^{2})$. In short, for every $i\in V_{1}$, we need to preserve at least $\Omega_{\varepsilon}(n)$ edges of the form $(i,j)$ with $j\in V_{2}$ in order to correctly estimate the total contribution of $\mathbf{R}(10)$ and $\mathbf{R}(11)$.
@@
[…]
For succinctness, we define $\vec{D}^{r}:=D_{1}\times\cdots\times D_{r}$. Given a valued relation $\mathbf{R}:\vec{D}^{r}\to\mathbb{Z}^{\geq 0}$ and a relation $R_{0}\subseteq\vec{D}^{r}$, we say that $\mathbf{R}\upharpoonright R_{0}$ is balanced if for all $t\in R_{0}$, $\mathbf{R}(t)$ has the same value. This motivates the definition of a generalized and $\and_{k}$.
@@
We say that an $r$-sorted valued relation $\mathbf{R}:\vec{D}^{r}\to\mathbb{Z}^{\geq 0}$ has a restriction to a generalized and $\and_{k}$ if there exists $E_{1}\subseteq D_{1},\ldots,E_{r}\subseteq D_{r}$ as well as a set of $k$ distinguished coordinates $S\subseteq[r]$ with the following properties.
@@
We say that $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a balanced and $\and_{k}$ if every nonzero output has the same value. Otherwise, we say that $\mathbf{R}\upharpoonright\vec{E}^{r}$ is an imbalanced and $\and_{k}$.
Note that $\mathbf{R}$ has a restriction to at least one generalized and $\and_{k}$ if and only if $\mathbf{R}$ has a restriction to and $\and_{k}$ (as in Definition 6.1).
@@
In this case, $\mathbf{R}\upharpoonright D^{2}$ is an imbalanced and $\and_{1}$ because the value of the first coordinate suffices to distinguish between zero and non-zero but the second coordinate still affects the value of the output. We can now state our classification.
@@
Let $C$ be the complete, $r$-partite instance of an $r$-sorted valued relation $\mathbf{R}:\vec{D}^{r}\to\mathbb{Z}^{\geq 0}$ with $n$ variables in each part, and let $c=\max\{k:\mathbf{R}\text{ has a restriction to }\mathbf{AND}_{k}\}$. If every generalized and $\and_{c}$ of $\mathbf{R}$ is balanced, let $\widehat{c}=c$. Otherwise, let $\widehat{c}=c+1$. Then:
@@ 6.3 Proof of the Lower Bound
If $\mathbf{R}$ has a restriction to and $\and_{k}$, then every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega(n^{k})$ constraints.
@@
If $\mathbf{R}$ has a restriction to an imbalanced and $\and_{k}$, then every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega(n^{k+1})$ constraints if $\varepsilon$ is a sufficiently small constant depending on the relation $\mathbf{R}$.
@@
Consider $E_{1}\subseteq D_{1},\ldots,E_{r}\subseteq D_{r}$ as well as a set of $k$ distinguished coordinates $S\subseteq[r]$ such that $\mathbf{R}\upharpoonright\vec{E}^{r}$ which is an imbalanced generalized and $\and_{k}$. Without loss of generality, we may assume that $S=[k]$ and $E_{1}=\cdots=E_{k}=\{0,1\}$ so that if $\mathbf{R}(t)>0$ for some $t\in\vec{E}^{r}$ then $(t_{1},\ldots,t_{k})=(1,\ldots,1)$.
Since $\mathbf{R}\upharpoonright\vec{E}^{r}$ is imbalanced, there exists $s,t\in\vec{E}^{r}$ with $0<\mathbf{R}(s)<\mathbf{R}(t)$. Since $\vec{E}^{r}$ is a product set, we can construct a chain of tuples $s^{0}=s,\ldots,s^{\ell}=t\in\vec{E}^{r}$ such that $\operatorname{Ham}(s^{i},s^{i+1})=1$ throughout the chain and $s^{i}_{j}=1$ whenever $j\in[k]$. Since $\mathbf{R}(s)\neq\mathbf{R}(t)$, there must be two consecutive terms in this sequence which differ in value. Thus, we may assume without loss of generality that $\operatorname{Ham}(s,t)=1$. Furthermore, we may assume that $s_{k+1}\neq t_{k+1}$ and $\mathbf{R}(s)<\mathbf{R}(t)$.
@@ 6.4 Proof of the Upper Bound
Since $\mathrm{sat}_{\mathbf{R},C}(\psi)>0$, there exists $e_{1}\in D_{1},\ldots,e_{r}\in D_{r}$ such that $(e_{1},\ldots,e_{r})\in\operatorname{supp}\mathbf{R}$ and $a_{i,e_{i}}\geq 1$ for all $i\in[r]$. Of all such choices of $(e_{1},\ldots,e_{r})$, pick the one which maximizes the number of $i\in[r]$ for which $e_{i}=d_{i}$. Now for all $i\in[r]$, let $E_{i}:=\{d_{i},e_{i}\}.$ By the maximality condition on the choice of $(e_{1},\ldots,e_{r})$, we have that $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a copy of and $\and_{k}$ for some $k\leq c$. Therefore,
@@
Moving forward, we assume that $\widehat{c}=c$. That is, every generalized and $\and_{c}$ appearing in $\mathbf{R}$ is balanced. With this assumption, we now proceed to prove a stronger upper bound of $\widetilde{O}_{\mathbf{R},\varepsilon}(n^{c})$ on the size of the sparsifier using the methods of Section 5. More precisely, let $R:=\operatorname{supp}\mathbf{R}$ and note that $c=\max\{k:R\text{ has a restriction to }\mathbf{AND}_{k}\}$. Therefore, we can use many results of Section 5 verbatim.
@@
By Proposition 5.6, we have that $t^{\prime}\in E_{1}\times\cdots\times E_{r}$ satisfies $t^{\prime}\in\operatorname{supp}\mathbf{R}$ if and only if $t^{\prime}_{i}=t_{i}$ whenever $t_{i}\neq 0$. Thus, $\mathbf{R}\upharpoonright\vec{E}^{r}$ is a generalized and $\and_{c}$ (Definition 6.3) since $\operatorname{Ham}(t)=c$. Since we assume that every generalized and $\and_{c}$ is balanced, we must have that $\mathbf{R}(s)=\mathbf{R}(t)$, a contradiction.
Second, assume for the sake of contradiction that $\mathbf{R}(t)<\mathbf{R}_{1}(t)$. Thus, there is $s\sim_{I}t$ such that $\mathbf{R}(s)>\mathbf{R}(t)$. In particular, $s\in\operatorname{supp}\mathbf{R}$. If $t\in\operatorname{supp}\mathbf{R}$ as well, then we can get a contradiction similar to the previous argument, so assume that $t\not\in\operatorname{supp}\mathbf{R}$. Note that $\operatorname{Ham}(s)\geq c+1$ as $s$ and $t$ can only differ in coordinates $i\in[r]$ for which $t_{i}=0$. Since $\mathbf{R}$ lacks and $\and_{c+1}$, there must be some $s^{\prime}\in\operatorname{supp}\mathbf{R}$ with $\operatorname{Ham}(s^{\prime})=c$ and for all $i\in[r]$ either $s^{\prime}_{i}=0$ or $s^{\prime}_{i}=s_{i}$. Now $s^{\prime}\neq t$ as $t\notin\operatorname{supp}\mathbf{R}$, and it follows that $\operatorname{supp}(s^{\prime})\neq\operatorname{supp}(t)$ as well. However, $\operatorname{supp}(s^{\prime})\cup\operatorname{supp}(t)\subseteq\operatorname{supp}(s)$. Therefore there exists an $i\in\operatorname{supp}(s)\setminus\operatorname{supp}(t)$ for which $(i,s^{\prime}_{i})=(i,s_{i})\not\in I$. This contradicts the facts that $s\sim_{I}t$. We can thus conclude $\mathbf{R}(t)=\mathbf{R}_{1}(t)$. ∎
@@
Proof of Theorem 6.4, Item 1.